# Edge-Only Universal Adversarial Attacks in Distributed Learning

### v1

@@ 6 Conclusion and Future Work
Experimental results demonstrated that our attacks can produce a strong adversarial effect, even when the accessible edge part consists of a few initial model layers only, causing a high number of mispredictions comparable to standard UAPs crafted with full knowledge of the model and output class distribution.
In our experiments, particular attention was also devoted to understanding the propagation of the crafted perturbations to the cloud part, revealing that controlling the unknown cloud decision process to predict certain target classes can be particularly challenging. Despite that, our investigations also revealed intriguing inter-class behaviors, which need to be better investigated in future work as discussed in Section 5. Another interesting direction for future research consists in exploring the impact of the proposed attacks on models augmented with defenses and investigate the development of new defense mechanisms in this distributed setting.
@@
Overall, we believe that the findings of this work will contribute to a broader understanding and awareness of the threats to which distributed learning is exposed.

### v2

@@ VII Discussion and Conclusion
Experimental results demonstrate that our attacks can induce strong adversarial effects, even when only the initial layers of the edge model are accessible. These result in a high number of mispredictions, comparable to standard UAPs crafted with full knowledge of the model and its output class distribution. We also investigated how the crafted perturbations propagate through the cloud component, highlighting the challenge of influencing an unknown decision-making process.
One important direction is the development of novel defense mechanisms specifically tailored to distributed inference settings, as discussed in Sec. VI-H. In particular, future work should address robustness at the interface between edge and cloud components. Another important direction is the extension of this attack paradigm to other domains, which is especially relevant for the IoT and sensing communities, including applications in deep signal processing.
@@
We believe that the findings of this work contribute to a broader understanding and awareness of the threats that distributed learning systems face.
Acknowledgements
This work was partially supported by project SERICS (PE00000014) under the MUR (Ministero dell’Università e della Ricerca) National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
Appendix A Split Point and Architectures
In the following, we provided additional details concerning the splitting points of the tested DNNs. Subsequently, we discussed the architecture used for defining the binary classifiers $g^{l}$ used for the edge-only attacks.
A-A Selection of the split points
Each of the considered neural networks (VGG16, ResNet50, WideResNet101, and MobileNetV2) is split under the assumption that their backbones can be schematically divided into four approximately equal depths. This hierarchical partitioning aligns with the internal structure of each model, which can be segmented into four sequential components (key layers $l_{1},l_{2},l_{3},$ and $l_{4}$, as defined in Sec. VI). This setup enables the analysis of different edge depths, ranging from $1/4$ up to the full backbone, in line with related work (e.g., [57]).
To clarify the specific layers used and ensure reproducibility of the experiments, we report the exact layer paths from the torchvision models in Table VI.
TABLE VI: Edge depths and corresponding PyTorch layers.
Depth MobileNetV2 VGG16 ResNet50 WRN101 ViT-B/S
1/4 features[4] features[9] layer1 layer1 blocks[1]
2/4 features[8] features[16] layer2 layer2 blocks[4]
3/4 features[12] features[23] layer3 layer3 blocks[6]
4/4 features[18] features[30] layer4 layer4 blocks[10]
A-B Architecture of the classifiers
The classifier used to distinguish between target features and other features adopts a simple architecture to prevent overfitting to a small set of targeted samples, enabling better generalization to a broader set of samples within the same target class. The pseudocode of its architecture is in Fig. 16.
For this purpose, the first layer is a convolutional module with a kernel size of 3 that takes as input the features from the selected layers. The input and output channels vary depending on the selected layers, as it corresponds to the channel size of the select edge layer $l$. The output of the convolution is then passed through an adaptive average pooling layer, allowing the model to focus on channel activation rather than spatial information. This design choice is essential to make the classifier agnostic to the position of the features (not particularly relevant in the case of universal perturbations) and focused solely on channel analysis. Finally, a fully connected layer with a hidden size of 128 is used.
The classifiers $g^{l}$ were trained using an Adam optimizer with default settings and a learning rate $\textit{lr}=1\times 10^{-3}$. The number of epochs was set to $50$, and the batch size (used for both training and attack optimization) was $100$. Please note that, since the number of target samples in $\mathcal{D}$ was generally lower than the number of non-target samples, we used weighted sampling to balance the probability of selecting target and non-target samples during training
Please note that, regarding the computational cost of the optimization, although this is generally not the primary concern in the UAP setting (since perturbations are assumed to be computed offline and then applied at runtime), the implementation of targeted attacks requires lightweight training on edge features, whereas the formulation of CosUAP attacks does not. In both cases, the optimization process is faster than in classic white-box UAP methods because backpropagation is performed only through a part of the model’s layers.
class g_Classifier($n$.Module):
def __init__(self, n_channels, num_classes=1):
super().__init__()
self.conv_block = $n$.Sequential(
$n$.Conv2d(n_channels, n_channels, 3, padding=1),
$n$.AdaptiveAvgPool2d(1),
$n$.Flatten()
) if num_channels > 1 else $n$.Identity()
self.fc_block = $n$.Sequential(
$n$.Linear(n_channels, 128),
$n$.ReLU(),
$n$.Linear(128, 1)
def forward(self, x):
return self.fc_block(self.conv_block(x))
Figure 16: Pseudocode in PyTorch for the binary classifier used to distinguish feature representations.